OpenStudy (anonymous):

p=2a+3b+4c solve for c

OpenStudy (anonymous):

p=2a+3b+4c subtract 2a p-2a=3b+4c subtract 3b p-2a-3b=4c divide by 4 (1/4)p-(1/2)a-(2/4)b=c